A man moves on a straight path.Hr travels total distance l.He travels half of the journey with speed 10ms.The remaining part of journey was covered with speed 20ms for half of the remaining time and with velocity 40ms for other half of remaining time.what is the average speed

Grade:11

1 Answers

sonika p
6 years ago
let the total distance be d
For the first half distance, 
time taken = distance/speed = (d/2) / 10 = d/20 …..(1)
For the second half distance, 
           first half time, speed =20m/s
           second half time, velocity =40m/s
distance covered = speed*time
                         d/2   = 20*t/2 + 40*t/2 = 30t
               d = 60t 
average speed =total distance/ total time
                          = d/(t1 + t) 
                             d/(d/20 + d/60) = 15m/s
Thus, average speed is 15m/s